MENTORING PAIRINGS

APPRENTICE Graduate students and e-mails

Name  @math.uchicago.edu
Mohammed Abouzaid (weeks 1-4)  mabouzai
Niles Johnson (weeks 3-8)  niles
Ben Lee (weeks 1-4)  Lee
Justin Noel  justin
Eric Patterson  eric
Wei Ren  wei

Organizer: Mohammed Abouzaid Weeks 5–8: Niles Johnson and Justin Noel Undergraduate students, e-mails, and mentors.

Name (18)  @uchicago.edu  mentor(s)
Padraic James Bartlett  padraic  Deepam Patel and Eric Patterson
Ryan Dietzman  gammon  Ben Lee and Wei Ren
Nabeel Ebeid  nabeel  Ben Lee and Wei Ren
Matthew Gonwa  gonwa  Mohammed Abouzaid, Justin Noel, Niles Johnson
Molly Hales  mkhales  Deepam Patel and Eric Patterson
Milan Kidd  mkidd  Mohammed Abouzaid, Justin Noel, Niles Johnson
Yee Fay Lim  yflim  Deepam Patel and Eric Patterson
Neil Lutz  njlutz  Deepam Patel and Eric Patterson
James Marrone  jmar  Ben Lee and Wei Ren
Jonathan Nieder  jrnieder  Ben Lee and Wei Ren
Pawan Patel  pawan  Mohammed Abouzaid, Justin Noel, Niles Johnson
Edward T. Ruggeri  ruggeri  Deepam Patel and Eric Patterson
Jenya Sapir  jmsapir  Deepam Patel and Eric Patterson
Matthew Tunnell  mtunnell  Mohammed Abouzaid, Justin Noel, Niles Johnson
Walter Ulmer  ulmer  Ben Lee and Wei Ren
Maya Vinokour  vinokour  Mohammed Abouzaid, Justin Noel, Niles Johnson
Muhammad Waliji  mhwaliji  Mohammed Abouzaid, Justin Noel, Niles Johnson
Jonathan Williams  jonathan  Ben Lee and Wei Ren

2 MENTORING PAIRINGS

YSP Graduate students and e-mails

Name  @math.uchicago.edu
David Balduzzi  balduzzi
Jeremy Copeland  copeland
Richard Cudney  cudney
Tullia Dymarz  tullia
Joseph Johns  johns
Sanjeevi Krishnan  sanjeevi
Chris Wallace  chris

Organizer: Jeremy Copeland Undergraduate students, e-mails, and mentors.

Name (22)  Program @uchicago.edu  mentor(s)
Sam Altschul  9-10  salt  Richard Cudney
Andrea Byrnes  7-8  aebyrnes  Jeremy Copeland and Tullia Dymarz
Amanda Comi  9-10  abc  David Balduzzi and Chris Wallace
Milind Desai  7-8  mdesai  Joe Johns and Sanjeevi Krishnan
Alexander Frankel  11-12  afrankel  Richard Cudney
Igor Gorodezky  11-12  igor2006  Jeremy Copeland and Tullia Dymarz
Ilya Grigoriev  9-10  ilyag  Joe Johns and Sanjeevi Krishnan
Moshe Hoffman  7-8  mozes  Jeremy Copeland and Tullia Dymarz
Martin Jankowiak  9-10  martinj  David Balduzzi and Chris Wallace
Evan Jenkins  11-12  ejenkins  Richard Cudney
Phillip Kalikman  9-10  plk  Joe Johns and Sanjeevi Krishnan
Brian Keto  7-8  bketo  Jeremy Copeland and Tullia Dymarz
Thomas Koberda  11-12  koberdat  Richard Cudney
Deniz Sehime Kusefoglu  7-8  deniz  Jeremy Copeland and Tullia Dymarz
Sara Lapan  9-10  slapan  Joe Johns and Sanjeevi Krishnan
Jeff Meyer  7-8  jsmeyer  Joe Johns and Sanjeevi Krishnan
Kyle Ormsby  11-12  kmo  David Balduzzi an Chris Wallace
James Pascaleff  11-12  jpascale  David Balduzzi and Chris Wallace
Allie Shapiro  7-8  allie  Jeremy Copeland and Tullia Dymarz
Weiwei Tan  9-10  cora7  David Balduzzi and Chris Wallace
Fang Zhao  7-8  roseanne  Joe Johns and Sanjeevi Krishnan
Kevin Choi  Sit in  kylchoi  David Balduzzi and Chris Wallace

MENTORING PAIRINGS 3

SESAME Graduate students and e-mails

Name  @math.uchicago.edu
Chris Bremer (weeks 5-8)  cbremer
Catalin Carstea  carstea
Matthew Day (weeks 1-4)  mattday
Nick Gurski  gurski
Courtney Morris  clmorris
Kate Ponto  kponto
David Rule  rule
Sergey Vasilyev  vasilyev

Organizer: Kate Ponto Undergraduate students, e-mails, and mentors.

Name (25)  @uchicago.edu  mentor(s)
Philip Ascher  pa7  Nick Gurski, Matthew Day (1-4), Chris Bremer (5-8)
William Cavendish  wpcavend  Kate Ponto
Trevor Davis  trevorld  Kate Ponto
Daniel Arthur Dickinson  dand  David Rule and Catalin Carstea
Jesse Farmer  farmerje  Nick Gurski, Matthew Day (1-4), Chris Bremer (5-8)
Stephen David Flood  sflood  David Rule and Catalin Carstea
William T. Gignac  wtgignac  Nick Gurski, Matthew Day (1-4), Chris Bremer (5-8)
Shawn Henry  shawnh  Courtney Morris and Sergei Vasilyev
Sayer Herin  sayerh  Nick Gurski, Matthew Day (1-4), Chris Bremer (5-8)
Josh Jones  jonesj  Courtney Morris and Sergei Vasilyev
David Levinson  levinson  David Rule and Catalin Carstea
Paul C. Mantz  pcmantz  Courtney Morris and Sergei Vasilyev
Jonathon Mason  jmason  Kate Ponto
Duff Neill  duff  Courtney Morris and Sergei Vasilyev
Joe Ochiltree  jjo  Kate Ponto
Annie Roberts  redoak84  Nick Gurski, Matthew Day (1-4), Chris Bremer (5-8)
Rodrigo Rodriguez  rrodrigo  David Rule and Catalin Carstea
Jeff Schonert  jeffrey1  Courtney Morris and Sergei Vasilyev
Jason Shapiro  shapiro  David Rule and Catalin Carstea
Rafal Sobota  rsobota  Nick Gurski, Matthew Day (1-4), Chris Bremer (5-8)
Sonya Swanson  sonja  David Rule and Catalin Carstea
Geoffrey Thompson  gzt  Nick Gurski, Matthew Day (1-4), Chris Bremer (5-8)
Alex V. Tomasik  tomasik  Courtney Morris and Sergei Vasilyev
Sokol Vako  vako  Courtney Morris and Sergei Vasilyev
Calder Wishne  calder  David Rule and Catalin Carstea